# Enabling Quick, Accurate Crowdsourced Annotation for Elevation-Aware Flood Extent Mapping

### v1

@@ 2 Design
Before implementing our solution, we conducted interviews with domain scientist collaborators (i.e., machine learning researchers building models for flood extent mapping) to understand the state-of-the-art and requirements for this project. From these discussions, we list the requirements for building an effective application for flood annotation for ML training data:
@@
  • R2: Annotation tools must be efficient and quick. Annotating a single flooded region usually involves labeling millions of pixels and multiple hours of work, and study areas commonly consist of many regions. Because of this, the acquisition of ground truth labels for training is extremely time-consuming. Tools that increase annotation productivity will have a large impact on improving ML researchers’ workflows.
  • R3: The annotation application must fully support the large data sizes of study areas. Previous annotation work[2] can only process small patches at a time, lowering productivity and leading to potential inaccuracies on patch borders. Performance on large data should be an important aspect of the new application design.
We also discussed our collaborators’ interest and concerns with crowdsourcing annotations to be used as training data. They expressed confidence that, while working on flood forecasting requires extensive experience and domain knowledge, data for flood detection (i.e. flood extent mapping) can accurately be labeled without expert knowledge by using aggregation of annotations from many participants. This is supported by previous work[42] which shows that aggregate annotation data seems to converge to the highest quality around 20-25 annotators for a particular study area, although this number likely changes depending on the difficulty of annotation region, background of participants, and tools being used. While optimistic about crowdsourcing, after discussion we were left with another requirement for our annotation system to support crowdsourcing properly:
@@
We implement our application to address these requirements in section 3. Feedback was continuously integrated into our application throughout the implementation process, as advised by best practices [56].
Figure 3: An example showing mesh creation with the heightmap meshing utility and rendering with aerial imagery RGB texture on a 1000 $\times$ 500 resolution example.
@@ 3 Implementation
FloodTrace consists of a web-based system that addresses the requirements gathered in section 2 for both crowdsourcing and direct annotation by researchers. To address R1, our system provides interactive 3D visualization of aerial imagery and DEM data (subsection 3.1). R1 and R2 are both addressed by providing efficient elevation-guided semi-automatic annotation tools (subsection 3.2). R4 is addressed with aggregate annotation visualization for viewing uncertainty and making corrections (subsection 3.3). For R3, our application is built using the GPU-accelerated WebGL[28] backend in Three.js[61]. In our experimentation, the largest data size we have tested is 12000 $\times$ 12000 pixels, which is much larger than flood extent mapping regions. We acknowledge that limitations are present in the case that available memory will not be sufficient to host the DEM data. We give a brief overview of our UI design in subsection 3.4.
A complete pipeline overview is given in Figure 2. Researchers first provide corresponding aerial imagery and DEM data to the server backend for computing the 3D mesh and topological data structures necessary for our methods. For crowdsourcing, these can easily be precomputed and served directly at a deployed frontend site, as in our user study ([anonymized link to demo]), removing the need for crowdworkers to be given raw data or access to the backend server. Once data is served to the frontend, annotation can be accomplished using elevation-guided annotation tools directly on an interactive 3D rendering. Once annotations are complete, they can be downloaded and, in the case of those created directly by researchers, used to train dependent ML models. In the case of crowdsourcing, annotations can be downloaded and submitted to researchers for aggregation. Researchers can then use these aggregated annotations with uncertainty visualization tools in the frontend to review and improve them before being used to train dependent ML models.
Figure 4: An example of simplified data and resulting segmentations at persistence thresholds $\epsilon$ = 0.02, 0.04, and 0.08 of the elevation data function range. Observing the differences in elevation data that come from topological simplification is quite difficult, but the effect is clear when looking at the varying contour tree segmentations produced.
@@ 3.1 Rendering
An important feature to address R1 is providing the user with an interactive 3D visualization to inform their annotation. This is important to ensure that pixels with high or low relative elevation do not get labeled incorrectly, labels for ambiguous pixels covered by clouds or tree canopy can be inferred, and seed pixels for the BFS tool can be selected effectively.
Upon receiving the input of RGB imagery data and DEM data, mesh creation is done in the backend as shown in Figure 2. We utilize the HMM heightmap meshing utility [19] in the backend to triangulate a mesh from the elevation image using the method of Garland and Heckbert [22]. The generated STL file is sent back to the web application, where it is visualized with the RGB imagery data texture mapped onto the mesh, as shown in Figure 3.
We acknowledge that it would be possible to render the height field directly with high accuracy in real-time [38, 15, 10], however in our initial testing, we found direct height field rendering in Three.js to impact the interactivity of our system for large datasets. Pre-computing a triangulated mesh allows for easy, interactive visualization on our front end. While triangulating a surface from DEM data has been shown to introduce more error than higher-order methods [39], we find this acceptable as the mesh itself is only used for exploration in our system, and selections for annotation by our tools are done on the full-resolution underlying data.
@@
To cater to the needs of our collaborators, we incorporated options that allow users to easily modify the mesh appearance and seamlessly switch between 3D and 2D views. Our collaborators expressed that having the ability to view the RGB imagery on a flat 2D mesh can be beneficial in certain situations, prompting us to include this feature. As users annotate a region, the pixels labeled as flooded or dry are overlaid with a translucent color mask of red or blue respectively. This feature can be toggled, and it is helpful to do so after annotating large features with semi-automatic tools. This allows the user to better focus on the satellite imagery underneath the annotation and ensure the area has been annotated correctly. These mesh appearance functionalities are implemented with custom WebGL fragment shaders for our Three.JS rendering.
Figure 5: An example of the selection of the same pixel as flooded with one click of the segmentation tool, first on the unsimplified elevation data and then on the data simplified with persistence threshold $\epsilon=0.16$. Initially, the tool labels almost no pixels, but when used at a higher simplification level it labels an entire lake feature.
@@ 3.2 Elevation-Guided Annotation Tools
Following R1 and R2, FloodTrace utilizes DEM data for elevation-guided annotation tools that are able to quickly create accurate, robust labels for large flooded regions. These tools are built for intuitive use by both domain experts and crowdworkers. We provide both extension-type and segmentation-type semi-automatic methods in the form of the BFS (subsubsection 3.2.1) and topological segmentation (subsubsection 3.2.2) tools respectively.
@@ 3.2.1 BFS Tool
This method of extension-type semi-automatic annotation was first proposed by Adhikari et. al [2], and we adapted it as a state-of-the-art tool in our application. This tool is used by annotating a seed pixel, which is then extended in all directions by connected downstream or upstream pixels (depending on whether the label is flooded or dry) in the DEM data. This accurately labels flooded or dry regions by taking advantage of the physical constraint that if a location is flooded, then its adjacent locations with a lower elevation will also most likely be flooded. In the same way, adjacent locations of dry areas that have higher elevations will most likely be dry. This approach accurately annotates regions along elevation border pixels, as the BFS stops where the elevation becomes higher or lower than the connected pixel, depending on if flooded or dry is being labeled. This approach also addresses the difficulty of annotating ambiguous pixels in the RGB imagery, such as those covered by tree canopy or clouds, by automatically labeling them using neighboring pixels.
Thanks to the 3D mesh rendering in FloodTrace, the user can purposefully select high-elevation flooded pixels or low-elevation dry pixels in order to label as large of a downstream or upstream area as possible, improving efficiency. Our application also provides a polygon BFS tool, where one can select points to form an arbitrary polygon, fill the polygon, and run a BFS selection from all points on the borders. This feature allows larger areas to be selected quickly while ensuring that the annotation’s border pixels correctly correspond to elevation changes.
@@ 3.2.2 Topology Segmentation Tool
We provide this tool as a novel method for semi-automatic annotation of flooded regions. We utilize TDA to create a segmentation-type tool that can quickly annotate regions according to features in the elevation data. We find that contour tree segmentations give us topological features in the DEM data that correspond well to landforms such as rivers and hills (Figure 1), with natural borders likely to correspond to flood levels.
As described in subsection 1.4, we utilize persistent homology to simplify our data at varying thresholds before computing segmentations on these, allowing for interactive selection of features at different levels of detail. To visualize the effectiveness of simplification by persistence for contour tree segmentation, we show an example region simplified at different persistence thresholds and the corresponding segmentations in Figure 4. By providing all these segmentations for annotation selection, users can quickly annotate huge areas of pixels that are obviously flooded or dry while also maintaining the ability to accurately label less persistent features.
@@
To compute segmentation and simplification in our system, FloodTrace utilizes the Python binding of the Topology Toolkit (TTK) [62], specifically its functions for simplification by persistence (which implement the works of Tierny and Pascucci[63] and Lukasczyk et al.[44]) and contour tree creation (which implement the work of Gueunet et al. [29, 30]). Given a DEM dataset, the backend computes contour tree segmentations to be used in the frontend for annotation Figure 2. Annotation is done simply and quickly with the segmentation tool by applying flooded or dry labels to the segment currently under the cursor. The user can quickly switch between different simplification levels in the UI, interactively changing the segmentations that will be used for selection to that persistence threshold. We show example tool usage in Figure 5, first on the raw elevation data of the region, then on the topologically simplified data at the persistence threshold 0.16 of the function range.
@@
Based on demo feedback from our collaborators, we provide two options for visualizing the segmentations of a dataset. The first option is to paint the borders of the current segmentations on the mesh in white. This is used to quickly see which simplification level leads to the desired level of detail by viewing how segmentation borders change when cycling through the levels. The second option is to highlight in red or blue (depending on flooded or dry potential label) the segment currently hovered over by pressing the highlight key. This is used to quickly check the coverage of the selected segment to confirm it contains fully flooded or dry imagery before labeling.
@@
To keep the interaction simple for the user study, we provided six simplification levels based on thresholds at logarithmic steps along the data’s normalized function range (i.e. $\epsilon$ = 0, 0.01, 0.02, 0.04, 0.08, 0.16). In the final application, though, users can choose the number of thresholds and persistence values themselves when submitting input data, and query the backend for more simplified data as needed. Users must switch between simplification levels in order to ensure that they do not label an area so large that it includes both flooded and dry regions or so little that they do not fully capture a flooded or dry feature, so using this tool does take some user experience and experimentation.
@@ 3.3 Aggregated Annotation Visualization
Before visualization, each annotation in a given set is transformed into a 2-dimensional array where red (flooded) pixels are made -1, transparent (unlabeled) pixels are made 0, and blue (dry) pixels are made 1. For our first view, the aggregate view, we find the mean for each pixel across annotations, giving a value in [-1, 1] where a score of -1 means all annotators agreed the pixel is flooded, 1 means all annotators agreed the pixel is dry, and scores closer to 0 mean the annotators either disagreed or chose not to label the pixel. These values are then visualized as a color-mapped texture on the 3D rendering within our application. We follow the standard of inversely mapping uncertainty to saturation in our colormap, but, because the underlying RGB imagery also contains important information for the researcher, we extend this by additionally mapping uncertainty to texture transparency. Pixels that annotators uniformly labeled as flooded or dry are colored with a more opaque red or blue, while this color becomes more white and transparent the more uncertain annotators were about its label. To simplify the texture being mapped, a sliding tool in our interaction menu gives users the ability to threshold pixels by their certainty, collapsing values to 0 if their absolute value is not greater than the threshold. As 0 corresponds to full transparency in our colormap, this effectively hides the value of pixels below the threshold. An example of this view aggregated from 45 unique annotations from our user study is shown in Figure 6 (A, B). This view was developed with collaborator feedback as a way to quickly understand the labels being used for training data and identify possible sources of error, helping to address R4. In addition, while dependent models can be trained using soft labels of flood and dry scores, the ability to explore certainty thresholds in our view allows researchers to find effective values for binarizing aggregate data into hard labels.
@@
In order to fully address R4, we create another view specifically for correcting aggregate annotations. This view, the variance view, highlights the areas where annotators disagreed the most, as these are likely to be inaccurately annotated. While the first view displays the mean pixel values for a set of annotations, this one displays the variance. We use a similar color mapping as before, with higher uncertainty (in this case variance) corresponding to whiter color, but because we want to focus the user on the areas of high uncertainty for correction, we inversely map uncertainty to texture transparency. This view is shown in Figure 6 (C). With this view, researchers can easily check the areas of high variance and correct them with our annotation tools if they find them obviously flooded or dry. The crowdsourced annotations can then be downloaded with the corrected areas labeled for more accurate model training. With these two views, we enable ML researchers to rely on crowdsourced annotation data by making the aggregated data understandable and easily fixable when errors occur.
Figure 7: UI for our application. Options and key reminders are blown up for visibility. Note the minimal screen space taken up by UI components so that the region being annotated can make use of the full screen.
@@ 3.4 User Interaction
The UI for FloodTrace is shown in Figure 7, with an interaction menu in the top right, toggles between erase/fill and flood/dry selection in the top left, key reminders in the bottom left, and checkpoint/finish buttons in the bottom right. The on-screen UI is limited in order to simplify user experience and allow the annotation mesh to take up the full view. Along with the semi-automatic tools, we also provide a brush to manually annotate if desired, such as to erase unwanted annotation.
Users make annotations by aiming the mouse cursor at an area of the mesh and pressing the chosen key of any of the annotation tools. Point BFS, segmentation, and brush tools immediately annotate the area underneath the cursor using their respective algorithms (with the brush simply painting a square) on key press. Polygon selection with BFS, however, occurs by pressing a key to repeatedly select desired points on the polygon to be drawn and then pressing another key to confirm the selection which then runs the BFS method to fill the pixels within that polygon and surroundings.
Annotation tools apply color to the annotation texture determined by whether one is annotating as dry (blue) or flooded (red). The current annotation type is toggled with a switch on the top left of the UI and it also affects whether the BFS method selects upstream or downstream pixels. Users can also toggle annotation mode from ’Fill’ to ’Erase’ in order to use all the different annotation tools to remove annotation color rather than paint on the texture. Undo and redo functionality are given to allow users to take back accidental or ill-thought-out annotations, and put them back again.

### v2

@@ 2 Design
Before creating our solution, we interviewed domain scientist collaborators (i.e., machine learning researchers building models for flood extent mapping) to understand the state-of-the-art and requirements for this project. Our collaborators were two junior researchers (PhD students) and one senior researcher (associate professor) who had each been working in the domain for multiple years. In the initial interview, we simply allowed these collaborators to voice their goals for an annotation tool without providing guidance, to not lead their conclusions. In this discussion, they highlighted the need for elevation-guided annotation, then focused on their usage of the state-of-the-art [2] and its limitations. From this interview, we created requirements R1 and R2. From then, we began developing our application while continuously integrating feedback, as advised by best practices [56]. To accomplish this, we conducted several unstructured interviews via in-progress meetings to demo the tool and request feedback. Towards the end of the development cycle, the application was also shared with the collaborators via a web deployment for them to experiment further. While much of the feedback received led to usability improvements, early progress meetings also led us to the creation of requirement R3, as the researchers emphasized the labor hours required for annotation. We list the complete requirements we developed for building an effective application for flood annotation for ML training data:
@@
  • R2: The annotation application must fully support the large data sizes of study areas. Previous annotation work[2] can only process small patches at a time, lowering productivity and leading to potential inaccuracies on patch borders.
  • R3: Annotation tools must be efficient and quick. Annotating a single flooded region usually involves labeling millions of pixels and multiple hours of work, and study areas commonly consist of many regions. Because of this, the acquisition of ground truth labels for training is extremely time-consuming. Tools that increase annotation productivity will have a large impact on improving ML researchers’ workflows.
Orthogonally, we also discussed our collaborators’ interest and concerns with crowdsourcing annotations to be used as training data. They expressed confidence that, while working on flood forecasting requires extensive experience and domain knowledge, data for flood detection (i.e. flood extent mapping) can accurately be labeled without expert knowledge by using aggregation of annotations from many participants. This is supported by previous work[42] which shows that aggregate annotation data seems to converge to the highest quality around 20-25 annotators for a particular study area, although this number likely changes depending on the difficulty of annotation region, background of participants, and tools being used. While our collaborators were optimistic about crowdsourcing, after discussion we were left with another requirement for our annotation system to support crowdsourcing properly:
@@
We implement our system to address these requirements in section 3.
Figure 2: An example showing mesh creation and rendering with aerial imagery RGB texture on a 1000 $\times$ 500 resolution example.
@@ 3 Implementation
FloodTrace consists of a web-based system that addresses the requirements gathered in section 2 for both crowdsourcing and direct annotation by researchers. To address R1, our system provides interactive 3D visualization of aerial imagery and DEM data (subsection 3.1). R1 and R3 are both addressed by providing efficient elevation-guided semi-automatic annotation tools (subsection 3.2). R4 is addressed with aggregate annotation visualization for viewing uncertainty and making corrections (subsection 3.3). For R2, our application is built using the GPU-accelerated WebGL[30] backend in Three.js[60]. In our experimentation, the largest data size we have tested is 12000 $\times$ 12000 pixels, which is much larger than flood extent mapping regions. We give a brief overview of our UI design in subsection 3.4.
A complete pipeline overview is given in Figure 1. Researchers first provide corresponding aerial imagery and elevation data to the server backend for computing the 3D mesh and topological data structures necessary for our methods. For crowdsourcing, these can easily be precomputed and served directly at a deployed frontend site, as in our user study, removing the need for crowdworkers to be given raw data or access to the backend server. Once data is served to the frontend, annotations can be made using elevation-guided tools on an interactive 3D rendering. Once the annotation process is complete, those can be downloaded and used directly to train ML models. In the case of crowdsourcing, annotations can be downloaded and submitted to researchers for aggregation. Researchers can then use these aggregated annotations with uncertainty visualization tools in the frontend to review and improve them before using them to train ML models.
@@ 3.1 Rendering
An important feature to address R1 is providing the user with an interactive 3D visualization to inform their annotation. Upon receiving the input of RGB imagery data and DEM data, our application creates a 3D mesh in the backend as shown in Figure 1. We utilize the HMM heightmap meshing utility [21] to triangulate a mesh from a given elevation image using the method of Garland and Heckbert [24]. The generated STL file is sent back to the web application, where it is visualized with the RGB imagery data texture mapped onto the mesh, as shown in Figure 2.
We acknowledge that it would be possible to render the height field directly [39, 17, 12], however in our initial testing, we found direct height field rendering in Three.js to impact the interactivity of our system for large datasets. Pre-computing a triangulated mesh allows for easy, interactive visualization on our front end. While triangulating a surface from DEM data has been shown to introduce more error than higher-order methods [40] and recent work has been able to reconstruct volumes for flooded terrains with higher accuracy [11, 8], we find this loss acceptable for our use case as the mesh itself is only used for exploration. Our elevation-guided annotation tools use the full-resolution underlying elevation data.
@@
Our collaborators expressed that having the ability to view the RGB imagery on a flat 2D mesh can be beneficial in certain situations, so we incorporate an option that allow users to seamlessly switch between 3D and 2D views. As users annotate a region, the pixels labeled as flooded or dry are overlaid with a translucent color mask of red or blue respectively. These colors were chosen to match the convention in the domain [2], with bright red and blue uncommon in satellite imagery. This overlay can be toggled, and it is helpful to do so after annotating large features with semi-automatic tools to ensure the area has been annotated correctly. These mesh appearance functionalities are implemented with custom WebGL fragment shaders for our Three.JS rendering.
Figure 3: An example region (A) with contour tree segmentations after no simplification (B) and at persistence thresholds $\epsilon$ = 0.02 (C), 0.04 (D), and 0.08 (E) of the elevation data function range. As seen in (B), segmentations produced without simplification are too noisy to be useful, while those in (E) correspond to data features such as the hills, lake, and river. Segmentations colored by rainbow colormap. Our application allows users to label these segmentations in one click to quickly annotate features as flooded or dry.
@@ 3.2 Elevation-Guided Annotation Tools
Following R1 and R3, FloodTrace utilizes elevation data for annotation tools that are able to quickly create accurate, robust labels for large flooded regions. These tools are built for intuitive use by both domain experts and crowdworkers. We provide both extension-type and segmentation-type semi-automatic methods in the form of the BFS (subsubsection 3.2.1) and topological segmentation (subsubsection 3.2.2) tools respectively.
@@ 3.2.1 BFS Tool
This method of extension-type semi-automatic annotation was first proposed by Adhikari et. al [2], and we adapted it as a state-of-the-art tool in our application. This tool is used by annotating a seed pixel, which is then extended in all directions by connected downstream or upstream pixels (depending on whether the label is flooded or dry) in the elevation data. This accurately labels flooded or dry regions by taking advantage of the physical constraint that if a location is flooded, then its adjacent locations with a lower elevation will also most likely be flooded. In the same way, adjacent locations of dry areas that have higher elevations will most likely be dry. This approach accurately annotates regions along elevation border pixels, as the BFS stops where the elevation becomes higher or lower than the connected pixel, depending on if flooded or dry is being labeled. This approach also addresses the difficulty of annotating ambiguous pixels in the RGB imagery, such as those covered by tree canopy or clouds, by automatically labeling them using neighboring pixels.
Thanks to the 3D mesh rendering in FloodTrace, the user can purposefully select high-elevation flooded pixels or low-elevation dry pixels in order to label as large of a downstream or upstream area as possible, improving efficiency. Our application also extends the method with a polygon BFS tool, where one can select points to form an arbitrary polygon, fill the polygon, and run a BFS selection from all points on the borders. This feature allows larger areas to be selected quickly while ensuring that the annotation’s border pixels correctly correspond to elevation changes.
@@ 3.2.2 Topology Segmentation Tool
We utilize topological data analysis to create a segmentation-type tool that can quickly annotate regions according to multi-scale features in the elevation data. To do this, FloodTrace utilizes the Python binding of the Topology Toolkit (TTK) [61], specifically its functions for simplification by persistence (which implement the works of Tierny and Pascucci[62] and Lukasczyk et al.[44]) and contour tree creation (which implement the work of Gueunet et al. [31, 32]). Given input data, the backend first uses persistent homology to compute simplifications of the data at varying thresholds, then creates a set of contour tree segmentations from the simplified data. These segmentations are sent to the frontend, where they can then be used for annotation. We show an example region with contour tree segmentations from data simplified at different levels in Figure 3. Notice how the simplified segmentations (such as E) capture useful terrain features for flooded and dry areas, such as the large hill at the bottom left, lake and river in the middle, and pond in the bottom right.
By providing the segmentation set for selection, users can quickly annotate large, obviously flooded and dry features, while maintaining the detail necessary to accurately label smaller features, allowing for robust annotation of features at different levels of detail. Annotation is done simply and quickly with the segmentation tool in our application by applying flooded or dry labels to the segment currently under the cursor. The user can quickly switch between different simplification levels in the UI, interactively changing the segmentations that will be used for selection to that level. Users must switch between simplification levels in order to ensure that they do not label an area so large that it includes both flooded and dry regions or so little that they do not fully capture a flooded or dry feature, so using this tool does take some experimentation. Example annotation using this tool can be seen in our supplementary video.
@@
Based on demo feedback from our collaborators, we provide two options for visualizing the segmentations of a dataset. The first option is to paint the borders of the current segmentations on the mesh in white, which effectively shows contours of the data simplified at the current level. This is used to quickly see which simplification level leads to the desired level of detail by viewing how segmentation borders change when cycling through the levels. The second option is to highlight in red or blue (depending on flooded or dry potential label) the segment currently hovered over by pressing the highlight key. This is used to quickly check the coverage of the selected segment to confirm it contains fully flooded or dry imagery before labeling.
@@
To keep the interaction simple for the user study, we provided six simplification levels based on thresholds at logarithmic steps along the datasets’ normalized function range (i.e. $\epsilon$ = 0, 0.01, 0.02, 0.04, 0.08, 0.16). While we found these sufficient for all of our study regions, necessary simplification could vary based on a particular researcher’s dataset. To support this, users can choose the number of thresholds and persistence values themselves when submitting input data, as well as query the backend for more simplified data as needed.
@@ 3.3 Aggregated Annotation Visualization
Before visualization, each annotation in a given set is transformed into a 2-dimensional array where red (flooded) pixels are made -1, transparent (unlabeled) pixels are made 0, and blue (dry) pixels are made 1. For our first view, the aggregate view, we find the mean for each pixel across annotations, giving a value in [-1, 1] where a score of -1 means all annotators agreed the pixel is flooded, 1 means all annotators agreed the pixel is dry, and scores closer to 0 mean the annotators either disagreed or chose not to label the pixel. These values are then visualized as a color-mapped texture on the 3D rendering within our application. We follow the standard of inversely mapping uncertainty to saturation in our colormap, but because the underlying RGB imagery also contains important information for the researcher, we extend this by additionally mapping uncertainty to texture transparency. Pixels that annotators uniformly labeled as flooded or dry are colored with a more opaque red or blue, while this color becomes more white and transparent the more uncertain annotators were about its label.
To simplify the texture being mapped, a sliding tool in our interaction menu gives users the ability to threshold pixels by their certainty. This collapses pixels; values to 0 if their absolute value is not greater than the threshold, effectively hiding the values since 0 is fully transparent. An example of this view aggregated from 45 unique annotations from our user study is shown in Figure 4 (A, B). This view was developed with collaborator feedback as a way to quickly understand the labels being used for training data and identify possible sources of error, helping to address R4. In addition, while dependent models can be trained using scored soft labels of flood and dry scores, the ability to explore certainty thresholds in our view allows researchers to find values for binarizing aggregate data into hard labels.
@@
In order to fully address R4, we create another view specifically for revising aggregated crowdsourced annotations. This view, the variance view, highlights the areas where annotators disagreed the most, as these are likely to need revision. It does this by displaying the computed variance for the set of annotations. We use a similar color mapping as before, with higher uncertainty (in this case variance) corresponding to white color. Using a white color for areas of interest can clash with the satellite imagery when it contains white features such as buildings, so we also provide the option to use a low saturation pink. Because we want to focus the user on the areas of high uncertainty for correction, we inversely map uncertainty to texture transparency. This view is shown in Figure 4 (C). With this view, researchers can easily check the areas of high variance and annotate them with our brush, BFS, or segmentation tools just as they would for a normal annotation task. The crowdsourced annotations can then be downloaded with hard correction labels applied to these annotated areas to improve model training.
Figure 5: UI for our application. Settings blown up for readability.
@@ 3.4 User Interaction
The UI for FloodTrace is shown in Figure 5, with an interaction menu in the top right, toggles between erase/fill and flood/dry selection in the top left, key reminders in the bottom left, and checkpoint/finish buttons in the bottom right. The on-screen UI is limited in order to simplify the user experience and allow the annotation mesh to take up the full view.
Users make annotations by aiming the mouse cursor at an area of the mesh and pressing the chosen key of any of the annotation tools. Along with the semi-automatic tools, we also provide a brush to manually paint and erase annotation. Point BFS, segmentation, and brush tools immediately annotate the area underneath the cursor using their respective algorithms on key press. Polygon selection with BFS occurs by pressing a key to repeatedly select desired points on the polygon to be drawn and then pressing another key to confirm the selection. This then fills the selected polygon and runs the BFS method on the pixels of its boundaries.
The current annotation type/color is toggled with a switch on the top left of the UI, which affects whether the BFS method selects upstream or downstream pixels. Users can also toggle annotation mode from ’Fill’ to ’Erase’ in order to use all the different annotation tools to remove annotation color rather than paint on the texture. Undo and redo functionality are given to allow users to take back accidental or ill-thought-out annotations, and put them back again.